A candle is lit and burns and the length of the candle changes at a constant rate of -1.9 inches per hour. 2 hours after the candle was lit the candle is 8.9 inches long.
a.Write a formula that expresses the remaining length of the candle in inches,
L, in terms of the number of hours t that have elapsed since the candle was lit.
b.Use your formula from part a. to determine how long the candle was before it was lit.

Answer by ikleyn(53751)

.
A candle is lit and burns and the length of the candle changes at a constant rate of -1.9 inches per hour.
2 hours after the candle was lit the candle is 8.9 inches long.

a.  Write a formula that expresses the remaining length of the candle in inches, 
L, 
    in terms of the number of hours t that have elapsed since the candle was lit.


        The remaining length  L(t) = 8.9 + 2*1.9 - 1.9*t  inches,   t >= 0.   (1)



b.  Use your formula from part a. to determine how long the candle was before it was lit.


        how long the candle was before it was lit = L(0) = simply substitute t= 0 into the formula (1) to get = 8.9 + 2*1.9 inches.